EVIDENCE OF PROTON TUNNELING IN THE ROTATIONAL SPECTRUM OF THE PROPIOLIC ACID-FORMIC ACID HYDROGEN BONDED COMPLEX

\underline{KEVIN O. DOUGLASS}, PAM L. CRUM, RICHARD D. SUENRAM, BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., P.O. Box 400319, Charlottesville, VA 22904.

We have measured the rotational spectrum of the hydrogen bonded system of propiolic acid and formic acid. The spectral fitting was aided by density functional calculations and FTMW-MW double resonance spectroscopy. The structure is predicted to be an 8 member ring system where the hydrogen bonding takes place between the acid functionalities of the monomers. The spectra were obtained using a neat 4:1 mixture of propiolic - formic acid while heating the nozzle to 50^{o}C. The spectrum was measured from 10 to 22 GHz. Each a-type rotational transition consists of a doublet with an intensity ratio that alternates with even and odd K_{a} and are suggestive of 3:1 hydrogen spin statistics. These doublets result from proton tunneling splitting. The results of the proton tunneling splitting will be presented.